OpenStudy (anonymous):

Simplify the expression using the properties of exponents: (x^-2y)^3(3x^4y^3)^2 ______________________ (4xy^-2)^2(x^2y^4)^2

OpenStudy (anonymous):

\[(x^{-2}y)^3(3x^4y^3)^2=x^{-6}y^39x^8y^6\] is the numerator. i multiplied the exponents. the denominator will be \[(4xy^{-2})^2(x^2y^4)^2=16x^2y^{-4}x^4y^8\] giving \[\frac{x^{-6}y^39x^8y^6}{16x^2y^{-4}x^4y^8}\]

OpenStudy (anonymous):

distributing the exponents outside the parenthesis:\[((x^{-6}y^3)*(9x^8y^6))\div((16x^2y^{-4})*(x^4y^8))\]

OpenStudy (anonymous):

then you can simplify in the numerator and denominator first, by adding exponents \[\frac{x^{-6}y^39x^8y^6}{16x^2y^{-4}x^4y^8}=\frac{9x^2y^9}{16x^6y^4}\]

OpenStudy (anonymous):

take it away satellite

OpenStudy (anonymous):

i have probably made a mistake by now. you can check it kellb \[\frac{9x^2y^9}{16x^6y^4}=\frac{9y^5}{16x^4}\]\]

OpenStudy (anonymous):

maybe? just a bunch of book keeping your math teacher gave to keep you out of trouble for a minute

OpenStudy (anonymous):

i know... its awful...